OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (-9, 6)? A. x- 3y = 9 B. x + 3y = 9 C. 3x - y = −9 D. 3x + y = 9

terenzreignz (terenzreignz):

Well, if you're in a hurry, you can plug in these points into the equations and see which equation works for both points.

terenzreignz (terenzreignz):

Otherwise, you can use these two points to determine the exact equation of the line~ with or without those choices.

OpenStudy (anonymous):

so i just plug in (3,2) and (-9,6) into x and y?

terenzreignz (terenzreignz):

Yes, that's one way to do it, arbitrary, but relatively quick... See which choice works for BOTH points.

OpenStudy (anonymous):

okay i just tottaly confused myself lol

terenzreignz (terenzreignz):

Okay, let's start with choice A... x - 3y = 9 and test the point (3,2) Replace x with 3 Replace y with 2 3 - 3(2) = 9 3 - 6 = 9 -3 = 9 Which is false, so it's not A.

OpenStudy (anonymous):

x + 3y = 9 and test the point (3,2) Replace x with 3 Replace y with 2 3 + 3(2) = 9 3+ 6 = 9 9=9 so choice b is ture? i think?

terenzreignz (terenzreignz):

Remember, BOTH points have to be tested... one is not enough...

OpenStudy (anonymous):

x+3y=9 -9+3(6)=9 -9+18=9 9=9

OpenStudy (anonymous):

so its b right?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Yup. Nice job :D

OpenStudy (anonymous):

yay! thank you for your time :D